# Thoughtbubbles: an Unsupervised Method for Parallel Thinking in Latent Space

### v1

@@ Abstract
22footnotetext: Currently at Google DeepMind.33footnotetext: Currently at OpenAI.
Current approaches for scaling inference-time compute in transformers rely on training them to emit explicit chain-of-thought tokens before producing an answer. While these methods are powerful, they are limited because they cannot be applied during pretraining and are limited to only serially-generated, natural-language verbalization to scale inference-time compute. In this work, we propose Thoughtbubbles, a transformer variant that natively performs parallel adaptive computation in latent space by learning to fork or delete residual streams. Thus, tokens that require a large amount of computation can form a “bubble” of cloned residuals in the middle of the network for additional thinking. Crucially, this behavior is learned during pretraining with only language modeling loss. Thoughtbubbles outperforms both standard decoder LMs as well as non-adaptive parallel computation approaches on OpenWebText and peS2o perplexity and in zero-shot evaluations such as HellaSwag and LAMBADA after pretraining across 150M to 772M parameter scales. The implicit nature of our method enables adaptive computation to be learned starting at pretraining time, paving the way to unify train and test-time behavior for reasoning models.
@@ 1 Introduction
Despite their unprecedented success, Transformers (Vaswani et al., 2017) only have a fixed computation budget and working memory, which present both a theoretical (Merrill & Sabharwal, 2023) and practical limit (Sanford et al., 2024) for solving complex, multi-step problems.
Due to the growing interest in extending the capabilities of transformers for difficult multi-step problems, many efforts are underway to surpass this bounded-computation limitation of transformers. The earliest and simplest is Chain of Thought (CoT) (Wei et al., 2023), where a transformer language model is explicitly prompted to provide a set of reasoning steps. This technique allows the model to break a problem down to subproblems, solve them individually, and cache intermediate results for the full solution—enabling a simple form of problem adaptivity (Merrill & Sabharwal, 2024).
Extending this result, Pfau et al. (2024) show both theoretically and practically that CoT improves the expressiveness of transformers—even when the CoT traces are replaced with a unique thinking token (dots) at test time: indicating that even residual streams alone, not serial recurrence, can improve computational performance.
Such an insertion of additional residual streams, the so-called “pause tokens”, has since become a growing trend of recent architecture research. Though methods vary in terms of where to actually insert the thinking tokens(Herel & Mikolov, 2024; Sun et al., 2025; Goyal et al., 2024), all pause token approaches insert additional computation streams prior to computation—limiting the model’s ability to allocate intermediate computation that is useful only in some, but not all layers (e.g., computation which is useful only after a few layers of attention). As Sun et al. (2025) notes, determining the location of these pause tokens often requires manual design following the structure of the problem, which may be intractable for general language models.
In response, we introduce Thoughtbubbles, a novel Transformer-based architecture which enables the unsupervised and dynamic allocation of additional parallel residual streams for extra computation and memory. We achieve this by introducing a novel forking mechanism between some layers, which computes and maintains a cumulative score for every residual stream and uses it to decide whether to create extra residuals or to delete existing ones.
@@
This formulation makes dynamic computation a budget-bounded allocation problem of these scores. In order to train these scores to be useful, we use these scores to mask both the model’s ability to attend to residual streams with low scores as well as limit the model’s ability to update them at each layer. This attenuation forces the model to provide higher scores to residual streams it deems more important, which will also result in increased forking of those streams. At the end of encoding, our model will decode each residual streams separately, including forked ones. Weighted by their scores, streams for the same token are then averaged together to produce the final distribution.
@@
We conduct a variety of pretraining experiments across 150M to 772M scales and make the following contributions:
@@
    We demonstrate that our approach performs better in validation perplexity as well in zero-shot evals of LAMBADA and HellaSwag against two competitive baselines—a regular parameter matched transformer, as well as a non-adaptive computation matched approach where the input residual is copied multiple times as filler tokens for additional computation before decoding. We additionally perform competitively against BLiMP and PIQA.
@@
    We further show that our method correctly allocates computation at interpretable regions of extra computation. In particular, our method allocates more computation at regions of higher uncertainty (i.e., posterior entropy).
We release a PyTorch implementation for the community. [^1] https://github.com/stanfordnlp/thoughtbubbles
@@ 2.1 Overall Architecture
Our architecture is a GPT-2-style decoder-only transformer (Radford et al., 2019), trained using the cross-entropy language modeling objective.
To achieve parallel computation, we want to allocate more residual streams corresponding to tokens that require more computation. To enable this, we propose a special type of transformer operation named “forking”, described in section 2.3, which can duplicate or remove some input residual streams for future computation.
The amount of forking is controlled by assigning a “cumulative score” between 0 and 1 to each residual stream. Each forking operation computes a “keep score” for each residual stream, which is multiplied to the cumulative score to update it, as well as a new “fork score” for the new residual. We describe the computation of these scores in section 2.3.
This setup reduces the dynamic computation task to determining which residual streams to keep or delete based on the value of cumulative scores: we take the top-k of the scores and perform their corresponding (i.e., keep / fork) actions. As long as the “useful” tokens receive the highest scores, the extra computation should help the performance of the model. To train the model to use the scores correctly, attention and residual updates are attenuated by the cumulative scores (section 2.4). That is, the tokens that the model needs to attend to and update the most become implicitly the highest-scoring tokens to be duplicated.
Additionally, we take special care about the RoPE position embeddings: we apply a “partial rotation” to the forked tokens proportional to the number of forks: the more forks a token has, the “closer together” each of their forks are. This design is described in detail in appendix D.
@@ 2.2 Notation
We will use $x_{j}^{(k)}\in\mathbb{R}^{d_{\text{model}}}$ to denote the $j^{\text{th}}$ residual stream at the $k^{\text{th}}$ layer. To emphasize that a particular token is the $j^{\text{th}}$ fork of token $i$, we will write $x_{i,j}^{(k)}$. We fork tokens to the left of the original input token. Thus, the original token is always $x_{i,0}^{(0)}$. A sequence of $q$ forks and original token can be written as $\quantity[x_{i,q}^{(k)}\dots x_{i,0}^{(k)}]$.
@@ 2.3 Forking
Given a set of scores for a layer $k$, we create a list $P=\quantity[\hat{p}_{\text{fork},0}^{(k)},\hat{p}_{\text{keep},0}^{{}^{\prime}(k)}\dots\hat{p}_{\text{fork},n}^{(k)},\hat{p}_{\text{keep},n}^{{}^{\prime}(k)}]$, we compute a top-k to downsample this list to obtain $P_{\kappa}$ where $|P_{\kappa}|=\kappa$. Using this list, we assemble the new residual stream set $X^{(k)}$ by the following two rules:
@@ 2.4 Residual Update Attenuation
$\text{Attn}\quantity(Q^{(k)},K^{(k)},V^{(k)})=\text{softmax}\quantity(\frac{Q^{(k)}{K^{(k)}}^{\top}+\mathbbm{1}\log\left(P^{(k)}\right)^{\top}}{\sqrt{d_{\text{model}}}})\quantity(V^{(k)}\odot P^{(k)})$ (8)
@@
$X^{(k)^{\prime}}$ $=\text{Attn}\quantity(f_{Q}\quantity(\text{LN}\quantity(X^{(k)})),f_{K}\quantity(\text{LN}\quantity(X^{(k)})),f_{V}\quantity(\text{LN}\quantity(X^{(k)})))\odot P^{(k)}\mathbbm{1}^{\top}+X^{(k)}$ (9)
@@
for $X^{(k)}$ being the cocatenated list of residual streams in the input of the layer, LN being layernorm, and $f_{Q,K,V}$ being the attention projections. If forking occurred prior to this layer, $X^{(k)}$ is as defined in eqs. 5 and eq. 6 , after forking takes place.
@@ 2.5 Output Averaging
$x_{i}^{(k)}=\frac{1}{\sum_{j}p_{\text{cum},\quantity(i,j)}^{(f)}}\sum_{j}p_{\text{cum},\quantity(i,j)}^{(f)}\text{softmax}\quantity(\text{Dec}_{\theta}\ \left(x_{i,j}^{(k)}\right)).$ (11)
@@ 2.6 Scoring and Sampling
Because of the possibility of varying $\kappa$ at inference time, there are two main ways inference can be performed in our model. Naively, we can set the inference budget $\kappa_{\text{inference}}$ to be the same as in training time $\kappa_{\text{train}}$, two or four times the block size atr training. We call this fixed forking. Alternatively, we can set the inference budget to be the same ratio as the training budget. $\kappa_{\text{inference}}$ is set to a value that maintains its same ratio to block size as during training; that, if $\kappa_{\text{train}}=2l_{\text{train}}$, then $\kappa_{\text{sample}}=2l_{\text{sample}}$. We call this dynamic forking, and discuss this method further in section E.1.
@@ 3.1 Parameter Selection and Training
Because our architecture takes token embeddings as input and produces token probabilities, it trains exactly like a regular decoder-only language model. As mentioned above, this means that the loss function can be standard language-modeling cross-entropy loss. Optimization is performed by the AdamW optimizer (Loshchilov & Hutter, 2017). Further optimization and architecture details can be found in appendix A.
We insert the first forking layer after a few regular transformer blocks to ensure that the forking score judgments see a broader context window. This is important in order to judge a token’s relative importance compared to the others. For all models in section 4, we train models at various scales with token forking placed prior to layers 3, 7, and 11. This means that for models with more layers, the majority of the latter half of transformer will contain no forking. We discuss this choice in appendix B.
3.2 Pretraining Datasets
We pretrain our approach on two datasets: OpenWebText (Gokaslan et al., 2019), a standard web-text pretraining corpus, as well as peS2o (Soldaini & Lo, 2023), a collection of academic papers sourced from the Semantic Scholar Open Research corpus (Lo et al., 2020). Pretraining is conducted for 2.5 billion tokens (75,000 steps). We sample batches at random positions throughout each dataset without additional masking.
@@ 3.4 Pretraining Evaluations
After pretraining, we conduct a variety of zero-shot evaluations on our models and baselines to examine their quality. They include the model’s measured perplexity on a holdout validation set, LAMBADA (Paperno et al., 2016) for context extraction, HellaSwag (Zellers et al., 2019) for common sense reasoning, BLiMP (Warstadt et al., 2020) for syntax understanding, and PIQA (Bisk et al., 2020) for embodied physical inference.
For each zero-shot downstream task, we use the dynamic budget as described in section E.1. We describe in detail the implementation of the zero-shot evaluations, including what to measure, in appendix E. Across all evaluations, we use the trained models as-is without additional fine-tuning.
@@ 4 Results
Figure 3: Dev-set perplexity of our approach and various baselines as a function of model scale on both OpenWebText and peS2o datasets. Across all scales, note that our method outperforms all baselines, including both computation and parameter-matched ones. Lower is better.
@@
Table 1: Zero-shot evaluation results across all model scales after pretraining on 2.5 billion tokens. Each setting is parameter-matched; baseline is a standard GPT-2-like model; copy-3 and copy-5 are models where the input residuals are copied multiple times and can attend to each other; ours is the thoughtbubbles transformer, with forking budget set to 2 $(\kappa=2L)$ and 4 $(\kappa=4L)$ times the input block size. The latter of which is roughly FLOPs-matched against copy-5 baseline.
@@
Across most zero-shot evaluations, our approach outperforms baselines.
For all LAMBADA and HellaSwag evaluations, we find that our approach outperforms both the parameter-matched baselines as well as the computation-matched baselines. However, we note that for BLiMP (syntax understanding) our model only outperforms the parameter-matched, but not computation-matched baselines—indicating that pruned dynamic parallel computation may not be as helpful for syntax matches. Finally, our model performs similarly to the baseline for embodied reasoning. We attribute this degraded performance to the fact that a short (2.5BT) training may not capture enough information for the embodied NLI to be effectively measured.

### v2

@@ Abstract
Current approaches for scaling inference-time compute in transformers train them to emit explicit chain-of-thought tokens before producing an answer. While these methods are powerful, they are limited because they cannot be applied during pretraining and rely solely on serially-generated, natural-language verbalization. In this work, we propose Thoughtbubbles, a transformer variant that natively performs parallel adaptive computation in latent space by learning to fork or delete residual streams. Thus, tokens requiring more computation can form a “bubble” of cloned residuals in the middle of the network. Crucially, this behavior is learned during pretraining with only language modeling loss. Using half of the training budget, Thoughtbubbles outperforms the perplexity and zero-shot evals of both standard decoder LMs and those using non-adaptive parallel computation approaches. These results hold across model sizes from 150M to 1.9B. Thoughtbubbles achieves competitive GSM8K results using half of the baseline’s token budget. The implicit nature of our method enables models to begin learning adaptive computation at pretraining time, paving the way to unified train-time and test-time scaling behaviors.
Neural Networks, Language Modeling, Adaptive Computation, Parallel Computation
@@ 1 Introduction
Despite their unprecedented success, Transformers (Vaswani et al., 2017) have a fixed computation budget and working memory, which present both a theoretical (Merrill and Sabharwal, 2023) and practical limit (Sanford et al., 2024) for solving complex, multi-step problems such as multi-hop retrieval or computer use agents.
Due to the growing interest in extending the capabilities of transformers for these difficult multi-step problems, many efforts are underway to surpass this bounded-computation limitation. The earliest and simplest is Chain of Thought (CoT) (Wei et al., 2023), which explicitly prompts the model to provide a set of reasoning steps. This technique allows the model to break a problem down into subproblems, solve them individually, and cache intermediate results for the full solution, thereby enabling a simple form of problem adaptivity (Merrill and Sabharwal, 2024).
Expanding upon this result, Pfau et al. (2024) shows both theoretically and practically that CoT improves transformers’ expressiveness. This result holds when CoT traces are replaced with a unique thinking token at test time, indicating that simply adding residual streams, without explicit token-level reasoning, can improve computational performance.
@@
Such an insertion of additional residual streams, the so-called “pause tokens,” has since become a growing trend in recent architecture research. Though methods vary in terms of where to actually insert the thinking tokens (Herel and Mikolov, 2024; Sun et al., 2025; Goyal et al., 2024), all pause token approaches insert additional computation streams prior to inference. These streams are also applied to all layers, limiting the model’s ability to allocate intermediate streams that are only useful in some layers (e.g., for computation that becomes useful after a few layers of attention have been applied). Furthermore, as Sun et al. (2025) notes, determining the location of pause tokens often requires manual design following the structure of the problem, which may be intractable for general language models.
In response, we present Thoughtbubbles, a novel Transformer-based architecture which enables the unsupervised and dynamic allocation of additional parallel residual streams for extra computation and memory. We achieve this by introducing a novel forking mechanism between some layers that computes and maintains a cumulative score for every residual stream and uses it to decide whether to create new residuals and keep existing ones.
This formulation makes dynamic computation a budget-bounded allocation problem of these scores. In order to train these scores to be useful, we use them to mask both the model’s ability to attend to residual streams with low scores as well as limit the model’s ability to update them at each layer. This attenuation forces the model to provide higher scores to residual streams it deems more important, which will also result in increased forking of those streams. At the end of encoding, our model will produce one output distribution for each stream by decoding each residual stream separately, including forked ones, and averaging the posterior probabilities weighted by their scores. We further find that a cheap approximation whereby the residuals themselves are averages still also confer similar expressivity advantages while giving a significant performance boost.
@@
We conduct a variety of pretraining experiments across 150M to 1.9B scales and make the following contributions:
@@
    We demonstrate that our approach performs better in all zero-shot evals as well as in GSM8K pass@1 using only half of the training token budget compared to baseline at 1.9B scale, and performance scales across 150M-1.9B scales
@@
    We further show that our method correctly allocates computation at interpretable regions of extra computation. Specifically, our method allocates more computation at regions of higher uncertainty (i.e., posterior entropy).
We release both PyTorch  [^1] https://github.com/stanfordnlp/thoughtbubbles and Jax  [^2] https://github.com/jemoka/fork-xla implementations for the community.
@@ 2.1 Overall Architecture
Our architecture is a decoder-only transformer (Radford et al., 2019), trained using the cross-entropy language modeling objective.
To achieve parallel computation, we want to allocate more residual streams corresponding to tokens that require more computation. To enable this, we propose a special type of operation named “forking”, described in Section 2.3, which can duplicate or remove some residual streams for future computation.
The amount of forking is controlled by assigning a “cumulative score” between 0 and 1 to each residual stream. This score can be interpreted as the stream’s existence indicator. In the forking operator, for each residual, the score is multiplied by two newly computed scores: “keep score” for updating the current stream, and a “fork score” indicating the importance of creating a new copy of this stream. We describe the computation of these scores in Section 2.3.
This setup reduces the dynamic computation task to determining which residual streams to keep or delete based on the value of cumulative scores: we take the top-k of the scores and keep their corresponding residuals (i.e., keep / fork). As long as the “useful” tokens receive the highest scores, the extra computation should help the performance of the model. To train the model to use the scores correctly, attention and residual updates are attenuated by the cumulative scores (Section 2.4). That is, the tokens that the model needs to attend to and update the most become implicitly the highest-scoring tokens to be duplicated.
Additionally, we take special care about the RoPE position embeddings: we apply a “partial rotation” to the forked tokens proportional to the number of forks: the more forks a token has, the “closer together” each of their forks are. This design is described in detail in Appendix F.
@@ 2.2 Notation
We will use $x_{j}^{(k)}\in\mathbb{R}^{d_{\text{model}}}$ to denote the $j^{\text{th}}$ residual stream at the $k^{\text{th}}$ layer. To emphasize that a particular token is the $j^{\text{th}}$ fork of token $i$, we will write $x_{i,j}^{(k)}$. We fork tokens to the left of the original input token. Thus, the original token is always $x_{i,0}^{(k)}$. A sequence of $q$ forks and original token can be written as $\quantity[x_{i,q}^{(k)}\dots x_{i,0}^{(k)}]$.
@@ 2.3 Forking
Given a set of scores for a layer $k$, we create a list $P=\quantity[\hat{p}_{\text{fork},0}^{(k)},\hat{p}_{\text{keep},0}^{{}^{\prime}(k)}\dots\hat{p}_{\text{fork},n}^{(k)},\hat{p}_{\text{keep},n}^{{}^{\prime}(k)}]$, we compute a top-k to shorten this list to obtain $P_{\kappa}$ where $|P_{\kappa}|=\kappa$. Using this list, we assemble the new residual stream set $X^{(k)}$ by the following two rules:
@@ 2.4 Residual Update Attenuation
$\tiny\begin{split}\text{Attn}\quantity(Q^{(k)},K^{(k)},V^{(k)})=\\ =\text{softmax}\quantity(\frac{Q^{(k)}{K^{(k)}}^{\top}+\mathbbm{1}\log\left(P^{(k)}\right)^{\top}}{\sqrt{d_{\text{model}}}})&\quantity(V^{(k)}\odot P^{(k)})\end{split}$ (8)
@@
$X^{(k)^{\prime}}$ $=\text{Attn}\left(f_{Q}(\text{LN}(X^{(k)})),f_{K}(\text{LN}(X^{(k)})),f_{V}(\text{LN}(X^{(k)}))\right)$ (9)
$\odot P^{(k)}\mathbbm{1}^{\top}+X^{(k)}$
@@
for $X^{(k)}$ being the cocatenated list of residual streams in the input of the layer, LN being layernorm, and $f_{Q,K,V}$ being the attention projections. If forking occurred prior to this layer, $X^{(k)}$ is as defined in eqs. 5 and Equation 6 , after forking takes place.
@@ 2.5 Output Averaging
$x_{i}^{(k)}=\sum_{j}\frac{p_{\text{cum},\quantity(i,j)}^{(f)}}{\sum_{l}p_{\text{cum},\quantity(i,l)}^{(f)}}\text{softmax}\quantity(\text{Dec}_{\theta}\ \left(x_{i,j}^{(k)}\right))$ (11)
@@
Although the above formulation has a principled probabilistic interpretation and also avoids the softmax bottleneck (Yang et al., 2018), it is expensive both computationally and memory-wise due to the large vocabulary sizes typically used by LLMs. Thus, for our most expensive, 1.9B parameter model, use the following cheap approximation instead:
$x_{i}^{(k)}=\text{softmax}\quantity(\text{Dec}_{\theta}\ \left(\sum_{j}\frac{p_{\text{cum},\quantity(i,j)}^{(f)}}{\sum_{l}p_{\text{cum},\quantity(i,l)}^{(f)}}x_{i,j}^{(k)}\right)).$ (12)
@@ 2.6 Scoring and Sampling
Because of the possibility of varying $\kappa$ at inference time, there are two main ways inference can be performed in our model. Naively, we can set the inference budget $\kappa_{\text{inference}}$ to be the same as in training time $\kappa_{\text{train}}$, two or four times the block size at training. We call this fixed forking. Alternatively, we can set the inference budget to be the same ratio as the training budget. $\kappa_{\text{inference}}$ is set to a value that maintains its same ratio to block size as during training; that, if $\kappa_{\text{train}}=2l_{\text{train}}$, then $\kappa_{\text{sample}}=2l_{\text{sample}}$. We call this dynamic forking, and discuss this method further in Section G.1. Note that dynamic forking is essential for keeping the forking distribution close to the training while doing autoregressive generation. Thus, because of its impracticality, the fixed forking approach is not used for any evaluations in our results; all autoregressive outputs are measured via dynamic forking.
@@ 3.1 Parameter Selection and Training
Because our architecture takes token embeddings as input and produces token probabilities, it trains exactly like a regular decoder-only language model. As mentioned above, this means that the loss function can be standard language-modeling cross-entropy loss. Optimization is performed by the AdamW optimizer (Loshchilov and Hutter, 2017) with more details described in Appendix A.
We insert the first forking layer after a few regular transformer blocks to ensure that the forking score judgments see a broader context window. This is important in order to judge a token’s relative importance compared to the others. For all models in Section 4, we train models at various scales with token forking placed prior to layers 3, 7, and 11. This means that for models with more layers, the majority of the latter half of transformer will contain no forking. We discuss this choice in Appendix D.
3.2 Training Recipe
1.9B Scale-Up Experiment
Our primary measurements are conducted using 1.9B parameter models trained by a standard two-stage training process. Our pre-training recipe, which we conduct for 40 billion tokens on a $10\%$ warmup-stable schedule followed by a constant learning rate. The training data is a mixture of $70\%$ FineWeb (Penedo et al., 2024), and $30\%$ peS2o (Soldaini and Lo, 2023) to create an English-dominant high-quality scientific text corpora. We then cool down these checkpoints using a cosine decay to $10\%$ of the original learning rate using a mid-training recipe, which involves a 2 billion token mixture of $5\%$ MMLU (Hendrycks et al., 2021), $30\%$ SmolTalk (Allal et al., 2025), $25\%$ of the pretraining mix, and finally $40\%$ GSM8K-aug—a synthetic math reasoning corpus widely used in small-scale reasoning literature (Ali et al., 2024; Shen et al., 2025; Kong et al., 2025). This forms a standard warmup-stable-decay schedule (Wen et al., 2025) used in current pretraining paradigms.
We checkpoint our model both at the end of pretraining as well as halfway through training, at 20B tokens—allowing the fully-trained baseline to be a generous over-estimate of both computation and data compared to our approach). The constant learning rate enables us to generate both of these checkpoints from the same training run.
150M-772M Scaling Suite
To demonstrate the scaling behavior of our approach, we pretrain our approach ranging from 150M - 772M scales on two datasets: OpenWebText (Gokaslan et al., 2019), a standard web-text pretraining corpus, as well as peS2o (Soldaini and Lo, 2023), a collection of academic papers sourced from the Semantic Scholar Open Research corpus (Lo et al., 2020). Pretraining is conducted for 2.5 billion tokens. These models use a more traditional cosine decay learning rate schedule.
@@ 3.4 Pretraining Evaluations
After pretraining, we conduct a variety of zero-shot evaluations on our models and baselines to examine their quality. They include the model’s measured perplexity on a holdout validation set, LAMBADA (Paperno et al., 2016) for context extraction, HellaSwag (Zellers et al., 2019) for common sense reasoning, BLiMP (Warstadt et al., 2020) for syntax understanding, AI2-ARC (Clark et al., 2018) for basic reasoning, and PIQA (Bisk et al., 2020) for embodied physical inference. For each zero-shot downstream task, we use the dynamic budget as described in Section G.1. We describe in detail the implementation of the zero-shot evaluations, including what to measure, in Appendix G. Across all evaluations, we use the trained models as-is without additional fine-tuning. For our full-scale model, after mid-training, we additionally evaluate our model and baselines’ zero-shot chain-of-thought performance on GSM8K’s test set (Cobbe et al., 2021).
@@ 4 Results
Ours ($\kappa=2L$) Baseline
20B tokens 40B tokens 40B tokens
Perplexity ($\downarrow$) 12.68 12.09 15.03
BLiMP ($\uparrow$) 78.54 78.41 77.49
HellaSwag ($\uparrow$) 50.04 54.16 47.29
PIQA ($\uparrow$) 73.42 73.00 71.79
ARC-Easy ($\uparrow$) 39.72 42.53 38.31
ARC-Challenge ($\uparrow$) 24.50 29.53 26.17
LAMBADA ($\uparrow$) 39.32 43.91 39.15
GSM8K ($\uparrow$) 31.50 32.10 31.46
Table 1: Zero-shot evaluation results across all model scales after pretraining on 20 and 40 billion tokens. Each setting is parameter-matched at 1.9B parameters. Baseline is a standard GPT-2-like model; ours is the thoughtbubbles transformer, with forking budget set to 2x $(\kappa=2L)$ the input block size. Measuring 20B tokens of pretraining of our approach against 40B tokens of baseline gives a generous temporal, data and computation lower bound.
@@
Table 2: Zero-shot evaluation results across all model scales after pretraining on 2.5 billion tokens. Each setting is parameter-matched; baseline is a standard GPT-2-like model; copy-3 and copy-5 are models where the input residuals are copied multiple times and can attend to each other; ours is the thoughtbubbles transformer, with forking budget set to 2x $(\kappa=2L)$ and 4x $(\kappa=4L)$ the input block size.
Figure 3: Dev-set perplexity of our approach and various baselines as a function of model scale on both OpenWebText and peS2o datasets. Across all scales, our method outperforms all baselines, including both computation and parameter-matched ones. Lower is better.
Even at half the token budget, at 1.9B scale, our model achieves better performance against all evaluation metrics.
Our model, as highlighted in Table 1, achieves better performance against baseline in all zero-shot evaluation metrics as well as on the reasoning dataset GSM8K. Excitingly, we reached or exceeded baseline performance by only using half of the token budget (i.e., half of the pretraining time), giving a generous lower-bound on our performance.
@@
Our approach’s performance dominates at scale.
Our approach’s performance dominates all baselines, and this effect becomes more prominent as model size scales. Across all scales, we find that our approach corfers a performance gain in all LAMBADA and HellaSwag evaluations in both the parameter-matched baselines as well as the computation-matched baselines. However, we note that for BLiMP (syntax understanding) our model only outperforms the parameter-matched, but not computation-matched baselines—indicating that dynamic parallel computation may not be as helpful for syntax matches.